[Halld-jlab] Hall D Staff Meeting Notes, August 25, 2014

Mark M. Ito marki at jlab.org
Tue Aug 26 15:22:21 EDT 2014


People,

Please find the notes below and at

https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_August_25,_2014 
.

   -- Mark
________________________________

Present:
Vladimir Berdnikov, Eugene Chudakov (chair), Mark Dalton, Alexandre 
Deur, Hovanes Egiyan, Sergey Furletov, Mark Ito, Vanik Kakoyan, Bert 
Manzlak, Lubomir Pentchev, Sergey Pozdniakov, Yi Qiang, Patrizia Rossi, 
Elton Smith, Sascha Somov, Simon Taylor, Tim Whitlatch, Beni Zihlmann

=Eugene=

* magnet testing last week a big success
** thanks to those who took shifts
** will run at 1200 A for commissioning
** magnet stays cold
** flow controller valve to be fixed
** ground fault came and went
** measuring the ground fault tripped the magnet twice
*** signals leading to trip confusing
* tomorrow: Accelerator Readiness Review (ARR)
** Eugene and Elton will present commissioning plan
** Alexandre, Eugene, Curtis, Elton developing materials
** active collimator should be responsive even with rough calibration
** question: what to do if main collimator way off
** Arne Freyberger's plan includes items from Hall D
** Schedule
*** start 9/29, beam up ramp, Hall closed, inset radiator, measure 
radiation, turn over beam to Hall D
*** with radiator measure position with profiler, take trouble-shooting data
*** one shift for trigger tuning
*** 50 nA
** 10/3 six day break, ACC on to other tasks
** 10/10 beam back to Hall D, tune trigger, take data, field off for 4 
days for FDC alignment
*** thin targets about 0.3% RL
*** will try several targets, positions
* Simon: simulation of commissioning geometry
** comm. target in
** beamline modification still to go in
* two progress reports left, then Director's Review, then Lehman Review
* Run Coordinator: Alexandre, Physics Division Liaison: Yi

=Tim=

* cleaning up last items: beam profile stand, start counter, microscope, 
beamline button up
* survey alignment scheduled, but falling behind schedule
* beam-pipes in tagger hall will be installed this week, thin window in
* tagger hall waiting on shielding blocks, arrive perhaps next week
** if all blocks do not come in, backup plan to do a single wall
* radiator foils being made, Anatoly and Slave working on them, wire also
** 1.5 micron, 10 micron, 30 micron aluminum foils

=Sascha=

* Thu. Nerses will check slow controls for bias voltages for pair 
spectrometer
* CTP story near end
** 15 CTPs have thermal pads replaced
** Fri. problem: could not talk to 2 boards, swapped out
** Hai: fans should go faster than 4k RPM
** Mark D.: F1-TDC not locking, need to increase fan speed
* controllers will go to tagger crates
** run with FCAL and BCAL[?]
* pair spectrometer course paddles connect to patch panel

=Sergey F.=

* DAQ hangs: testing nodes from computer center, estimating performance
** testing different configurations
* can run at 8 kHz now, bottleneck not clear
* preparing for monitoring

=Sergey P.=

* all terminal connection established
** cables installed
** not all terminal servers DHCP'ing
* lack of 15 pig tails

=Lubomir=

* chiller, thermostat adjusted\
** technician installing heat exchanger
* data with and without magnetic field, no surprises
** 1 wire missing, a TDC channel
** no change during fast dump
* commissioning
** software for alignment almost ready for all packages
** independent package, decodes EVIO, has translation table, tracking, 
analysis, 3-D viewer

=Beni=

* magnet-on running with cosmic trigger done
** 24 hours with magnet on
** CDC looks the same after a quick look
** curved tracks seen
* started using alcohol bubbler
* data taking during the weekend: TS crate died, fan control system went out
** take data again tomorrow
* talked to Tom Carstens about re-installing horizontal cosmic trigger 
paddles

=Mark=



=Simon=

* working on commissioning target
* updating start counter geometry
* moved recalculation of mass hypothesis from analysis to reconstruction
* tracked down memory leak in hdgeant
* translation tables checked for FDC
* work proceeding on CDC alignment, including straight line tracking for 
upcoming run

=Yi=

* boring running with solenoid (boring = good); thanks to shift takers
** no reboot needed
** alarm system worked
** want to automate ramping procedure
** refrigerator did well; might allow running at 1400 A
* Eugene: 3 questions from management
*# What was the cause of the quench?
*# What does GlueX really need?
*# Why not run at higher field?
** George Biallas organizing committee to look at these questions
* micro-channel plate sample sent to FNAL for beam tests, will be sent to us

=Elton=

* NIM electronics for FCAL: want to replace with VME if funds available
* revisiting Interface Control Document with Accelerator (ACC)
** Jim Stewart drafted it
** outlines responsibilities for systems along the beamline, specifies 
beam parameters

=Hovanes=

* ERR asked for controls feature for collimator: prevent extraction
** Vanik implemented change, tested it
* second beam profiler, use twin brother from Hall B
** read-out electronics: Struck scalars
** will postpone porting drivers to Linux, use VxWorks
** will need portable VME crate
* pair spectrometer
** voltage controls going in
* CAEN disconnect problem
** Nerses found way in driver to detect disconnect, can now make an alarm
* Vanik discovered bugs in CSS, group operations get stuck
** will try to catch exception and re-try
** will test tomorrow during DAQ tests

=Alexandre=

* looking at stability of PS field
* radiation monitoring in Hall and tagger few cables needed, RadCon will 
make them
* will meet with RadCon and Hovanes on controls
* last week started attending ACC meetings

=Mark D.=

* TDCs: some not working, not locking
** 5 channels dead, 10 chips on south = 40 channels
** can fix with changing bin from 56 to 58 ps
** getting 3 cm RMS for LED light pulses

=Bert=

* site office interested in observing fall protection practices
** work examples sought for observation
* ARR review preparations in progress

=Elton: SAF113 Training: Hall D Walk-Through=

* re-certification will be required



-- 
Mark M. Ito, Jefferson Lab, marki at jlab.org, (757)269-5295



More information about the Halld-jlab mailing list